[Brahms-l] Shift report 20050314 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Mon Mar 14 2005 - 08:00:00 EST
Shifter leader was kh
Shifters were 


After solving the machine problems left over from the previous shift, we finally got beam at about 3:00am.  It has been smooth running since then.
Runs accomplished:
Run   tr3   tr6
13776 82k   214k
13777 56k   162k
13778 50k   151k
13779 60k   190k

That brings the total for MRS 90 deg 700B to 838k tr2 out of a goal of 1M and for FS 6 deg 861A to 1.66M tr6 out of a goal of 2M.

Beam is currently being dumped.


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Mar 14 08:00:38 2005

This archive was generated by hypermail 2.1.8 : Mon Mar 14 2005 - 08:00:48 EST